OpenStudy (kyrie_hitsugia):

Find the value of the following expression: (38 ⋅ 2−5 ⋅ 90)−2 ⋅ 2 to the power of negative 2 over 3 to the power of 3, whole to the power of 4 ⋅ 328 (5 points)

OpenStudy (hoot_hoot):

\[[[[(38 \times 2 - 5 \times 90) - 2 \times 2]^2] \div 3^3]^4 \times 328\] Do the equation first inside the parentheses. \[= (38 \times 2) - (5 \times 90)\] When you get the answer, subtract it to \[(2 \times 2)\] because of MDAS (multiplication, division, addition, and subtraction). When you got the answer, square it. Then, you will divide it by 27 because of \[3^3\], and then solve for the answer raise to four: \[[(38 \times 2 - 5 \times 90)^2 \div 27]^4\]. When you get the anser, multiply it by 328. Hope you get the answer. :)